OpenStudy (anonymous):

@haleyeliabeth2017 @SolomonZelman let me post it may take a minute

OpenStudy (anonymous):

Which would not help to calculate the volume of the composite figure shown? \[lwh\] \[\pi dh\] \[\frac{ 1 }{ 3 }S ^{2}h\] \[\frac{ 1 }{ 3 }lwh\]

OpenStudy (haleyelizabeth2017):

@SolomonZelman